according to nscb statistics, the life expectancy of filipino women is 70.1 years. suppose a random sample of 30 women in town b yields a mean of 72.5 years with a population variance of 4.76 years, would you say that the life expectancy of the women in town b is greater than the average


a. Formulate the null and alternative hypothesis in sentence and in symbol

b. Determine what test statistic to use

c. Compute the test statistic


1
Expert's answer
2021-06-24T18:44:20-0400

a.

The null hypothesis "(H_0)" is: the life expectancy of the women in town b is not greater than the average.

The alternativel hypothesis "(H_1)" is: the life expectancy of the women in town b is greater than the average.

The following null and alternative hypotheses need to be tested:

"H_0:\mu\leq 70.1"

"H_1:\mu>70.1"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.


Based on the information provided, the significance level is "\alpha=0.05, df=n-1=30-1"

"=29" degrees of freedom, and the critical value for a right-tailed test is "t_c=1.699127."

The rejection region for this right-tailed test is "R=\{t:t>1.699127\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{72.5-70.1}{4.76/\sqrt{30}}\approx2.7616"

Since it is observed that "t=2.7616>1.699127=t_c," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is greater than "70.1", at the "\alpha=0.05"


Using the P-value approach: The p-value for right-tailed, "\alpha=0.05, df=29, t=2.7616" is "p=0.004937," and since "p=0.004937<0.05=\alpha," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is greater than "70.1", at the "\alpha=0.05"


Therefore, there is enough evidence to say that the life expectancy of the women in town b is greater than the average, at the "\alpha=0.05"